OpenStudy (tomektomeknyc):

If tan 3 theta = 1, 0 < theta < Pi , then which of the following is true: a) theta = arctan 1/3 b) theta = pi/12 c) theta = 1/3 arc tan pi/4 d) sec 3 theta = 2

OpenStudy (loser66):

your idea?

OpenStudy (tomektomeknyc):

not sure

OpenStudy (loser66):

try once, please.

OpenStudy (loser66):

\(tan(3\theta) =1\) \(3\theta =??\)

OpenStudy (tomektomeknyc):

1/tan

OpenStudy (loser66):

not that, it is \(3\theta = arctan 1\) are you familiar with it?

OpenStudy (tomektomeknyc):

not yet

OpenStudy (loser66):

we can go back and forth between an angle and its trigs, For example \[sin\dfrac{\pi}{4}=\dfrac{\sqrt{2}}{2}\\\rightarrow \dfrac{\pi}{4}=arcsin(\dfrac{\sqrt{2}}{2})\] your case is tan \[tan(3\theta)=1\\\rightarrow 3\theta = arctan(1)\]

OpenStudy (loser66):

you can go another way like this, tan (??) = 1 and we know, tan (pi/4) =1